OpenStudy (anonymous):

-4=2/5 *x a.10 b.-10 c.-1 d.1

OpenStudy (anonymous):

first change the problem around so it looks like -4 = 2x/5 then you mutliply 5 on both sides 5 x (-4) = 2x/5 x 5

OpenStudy (anonymous):

5 x (-4) = 2x what is 5 x (-4)?

OpenStudy (anonymous):

5 x (-4) = -20 so now we have -20 = 2x

OpenStudy (anonymous):

now we just need to dide 2 on both sides -20/2 = 2x/2 -20/2 = x

OpenStudy (anonymous):

so what is -20/2?

OpenStudy (anonymous):

-10 thank you so much i have a 2 more i need help on

OpenStudy (anonymous):

okay and no problem

OpenStudy (anonymous):

which equation is an identity? a.8-(5x+2)=-5x-6 b.7z+10-z=8z-2(z-5) c.8m-4=5m+8-m d.6y+5=6y-5 which equation has no solution? a.7v+2=8v-3 b.3x-5=3x+8-3 c.4y+5=4y-6 d.7z+6=-7z-5

OpenStudy (anonymous):

learn abou identities here - http://www.mathsisfun.com/definitions/identity.html not sure about the second one

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

no problem